OpenStudy (anonymous):

differentiate f(x) = xe^2cscx

OpenStudy (turingtest):

is this \[f(x)=xe^{2\csc x}\]?

OpenStudy (anonymous):

i mean e^x

OpenStudy (anonymous):

xe^xcscx

OpenStudy (turingtest):

\[f(x)=xe^x\csc x\]?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

well you know the product rule, right?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

so split this sucker up into two parts, like\[f(x)=(xe^x)\cdot(\csc x)\]then the derivative becomes\[f'(x)=(xe^x)'(\csc x)+(xe^x)(csc x)'\]

OpenStudy (anonymous):

\[e^xcscx+xe^x-cscxcotx\]?

OpenStudy (turingtest):

maybe i didn't do it lol

OpenStudy (turingtest):

hm no seems a bit off do you follow my explanation?

OpenStudy (anonymous):

yeah

OpenStudy (turingtest):

i think you just made a small error somewhere. your answer is very close, just missing a csc term

OpenStudy (anonymous):

i don't get it

zepdrix (zepdrix):

Product rule works the same way with three terms :d \[\Large\rm (uvw)'=u'vw+uv'w+uvw'\] \[\large\rm (x e^x \csc x)'=(x)' e^x \csc x+x(e^x)' \csc x+ x e^x(\csc x)'\]

OpenStudy (anonymous):

thanks

zepdrix (zepdrix):

So your middle term just needs to be fixed a little bit :)